OpenStudy (anonymous):

What is the solution of \[\log _{2x+3} 125 = 3\]

OpenStudy (anonymous):

x=1/3 x=1 x=7/3 x=4

OpenStudy (swissgirl):

Lets rewrite it as \( (2x+3)^3=125 \)

OpenStudy (anonymous):

2\[2x^{3} + 27 = 125\]

OpenStudy (swissgirl):

nope \((2x+3)^3 = (2x+3)(2x+3)(2x+3) \)

OpenStudy (anonymous):

Do you multiply from there?

OpenStudy (swissgirl):

hmm I am thinking of taking an easier route

OpenStudy (swissgirl):

Let w=2x+3 so we can rewite it as \( w^3=125\) did u follow that?

OpenStudy (anonymous):

what would do after

OpenStudy (swissgirl):

u didnt follow lol ok ill just continue maybe ull follow what i did once u see the full picture

OpenStudy (anonymous):

probably lol

OpenStudy (swissgirl):

\( \large \sqrt[3]{w^3 }= \sqrt[3]{125} \) \( \large w= \sqrt[3]{125} \) \(w=5\)

OpenStudy (swissgirl):

I basically tried to isolate the w or in other words solve for w

OpenStudy (anonymous):

Sadly, thats not one of the answers

OpenStudy (swissgirl):

now we know that w=2x+3 so lets subsitute our w for 2x+3 2x+3=5 2x=2 x=1

OpenStudy (swissgirl):

hahahah that is your answer lol

OpenStudy (anonymous):

1 is an answer

OpenStudy (swissgirl):

U so didnt follow what i did but alright at least u got the answer :)

OpenStudy (anonymous):

thanks lol

OpenStudy (anonymous):

What is the solution to the equation \[\frac{ 1 }{\sqrt{8}} = 4^{(m+3)}\]

OpenStudy (anonymous):

x=-15/4 x=-11/4 x=5/4 x=9/4